the frequency of revolution of electron in nth bohr orbit is Vn. Then what is the graph between log n and log (Vn/V1) ?

Grade:12th pass

1 Answers

Vikas TU
5 years ago
frequency is proporrional to the z^2/n^3
where z is the atomic nuber and n is the quantum number respectively.
Vn = k*(z^2/n^3)...................(1)
V1 = k*(z^2)...........................(2)
Dividing both eqns. we get,
Vn/V1 = 1/n^3
Taking log both sides we get,
log(Vn/V1) = -3log(n)
It can be written in a stratight line between log n and log (Vn/V1) as =>
x = -3y
or
y = -x/3 which is a straaight line with negative slope and intercept zero which means it will pass thorugh origin.